OpenStudy (anonymous):

~*~*~*~HELP~*~*~*~ Rational expression 3/x+1 - 3/x

sam (.sam.):

Common denominators

OpenStudy (anonymous):

how would u do that

OpenStudy (anonymous):

la la i love kpop.....okay buh bye

sam (.sam.):

Multiplying each other's denominators \[(\frac{3}{\color{red}{x+1}}\times\color{blue}{\frac{x}{x}})-(\frac{3}{\color{blue}{x}}\times\color{red}{\frac{x+1}{x+1}})\]

OpenStudy (anonymous):

\[\frac{3}{x+1} - \frac{3}{x} =\frac{3x -3(x+1)}{x(x+1)} =\frac{3x -3x-3}{x(x+1)1} \] \[=\frac{-3}{x(x+1)} \]

sam (.sam.):

Do you understand?

OpenStudy (anonymous):

yes thank you guys!(: